OpenStudy (anonymous):

someone want to help me multiply and divide complex numbers in polar form? :P thanks :) ill give a medal and fan ^_^

jimthompson5910 (jim_thompson5910):

go ahead and post the problem

OpenStudy (anonymous):

given z=6(cos 108deg+isin108deg) and w=7(cos 26deg+isin26deg), find and simplify z/w. round numerical entries in the answer to two decimal places.

OpenStudy (anonymous):

step by step how to please?

jimthompson5910 (jim_thompson5910):

It turns out that there is a very handy formula when doing this in polar form if x = r1*cis(theta1) y = r2*cis(theta2) Then x/y = (r1/r2)*cis(theta1 - theta2)

jimthompson5910 (jim_thompson5910):

cis(theta) is shorthand for cos(theta) + i*sin(theta)

jimthompson5910 (jim_thompson5910):

In this case, z = 6*cis(108) w = 7*cis(26)

jimthompson5910 (jim_thompson5910):

So, z/w = (6/7)*cis(108-26) z/w = (6/7)*cis(82) z/w = (6/7)*[cos(82) + i*sin(82)]

jimthompson5910 (jim_thompson5910):

now use a calculator: cos(82) = 0.13917310096007 (approximately) sin(82) = 0.99026806874158 which means we can now say z/w = (6/7)*[cos(82) + i*sin(82)] z/w = (6/7)*[ 0.13917310096007+ i*0.99026806874158] z/w = (6/7)*0.13917310096007+ (6/7)*i*0.99026806874158 z/w = 0.11929122939434 + 0.8488012017785i z/w = 0.12 + 0.85i

OpenStudy (anonymous):

thank you so much ^_^ do you know the formula for multiplying stuff like that too?

jimthompson5910 (jim_thompson5910):

if x = r1*cis(theta1) y = r2*cis(theta2) Then x*y = (r1*r2)*cis(theta1 + theta2) ----------------------------------------- if x = r1*cis(theta1) y = r2*cis(theta2) Then x/y = (r1/r2)*cis(theta1 - theta2)

jimthompson5910 (jim_thompson5910):

There are the 2 (very related) rules together

OpenStudy (anonymous):

thank you so so so so so so much!!!! :D

jimthompson5910 (jim_thompson5910):

you're welcome